OpenStudy (anonymous):

A man walks up 130 steps in 17 seconds. Considering each step is 20 cm and the person's weight is 196 lbs, how much mechanical power was done & how much mechanical power generated?

OpenStudy (anonymous):

First calculate the work done by gravity. Note it's negative, but will become positive taking direction of acceleration due to gravity into account. $$W_g=-hgm=-((130\,steps) \cdot (0.20\,m)\cdot(-9.8m/s^2)\cdot\frac{196\,lb}{2.2\,lb/kg})$$ Then divide by time to get the work done per second (power) \[P=\frac{W_g}{17\,s}\]

OpenStudy (anonymous):

Okay, so is mechanical work 22,700.36? or 254.8? or neither? And how much mechanical power was generated?

OpenStudy (anonymous):

Simply work out P. This will be the mechanical power generated. Remember power is ALWAYS i units joules/second. This is howmuch energy is expended per second. Then you could ask the question like, "if he only walks up for 2 seconds how much energy does he use?" then you would multiply you power (joules per second) to get how many joules he spends.